OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -4 + 5 + 14 + 23 + ... + 131

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

so -4, + 5, + 14, + 23, + ... + 131 so it starts off at -4 ends at +131 from -4 to +5 is 9 units so that'd be our "d" "common difference" since is an "arithmetic" sequence so.... what's 131? well.. .we dunno but we know -4 and "d" so \(\bf a_{\color{brown}{ n}}=a_1+({\color{brown}{ n}}-1){\color{blue}{ d}}\implies a_{\color{brown}{ n}}=-4+({\color{brown}{ n}}-1){\color{blue}{ 9}} \\ \quad \\ \textit{so what's term is it with 131?} \\ \quad \\ 131=-4+({\color{brown}{ n}}-1){\color{blue}{ 9}}\implies 131=-4+9n-9\implies 131=-13+9n \\ \quad \\ 131+13=9n\implies \cfrac{\cancel{ 144 }}{\cancel{ 9 }}=n\)

OpenStudy (anonymous):

n=16

OpenStudy (jdoe0001):